You have decided to place $166 in equal deposits every month at the beginning of the month into a savings account earning 5.58 percent per year, compounded monthly for the next 15 years. The first deposit is made today. How much money will be in the account at the end of that time period?

Please provide steps and show work using excel.
Answer #1

A 2 Deposit Amount (PMT) 3 Rate 4 NPER $166 0.4650% Rate per month 180 Months 6 FV $46,802.02в 1 2 Deposit Amount (PMT) 3 Rate 4. NPER 166 =5.58%/12 =15*12 Rate per month Months 6 FV =FV(B3,B4,-B2,,1)